# Conformal Prediction for Long-Tailed Classification

### v2

@@ 2.2 Approach II: Softening Classwise CP via Interp-Q
We now propose a simple way to interpolate between the behaviors of Classwise and Standard by linearly interpolating their quantile thresholds. We call this procedure Interp-Q for “interpolated quantile” because it constructs sets as ${\mathcal{C}}_{\textsc{Interp-Q}}(X):={\mathcal{C}}(X;\hat{\mathbf{q}}_{\textsc{IQ}})$ where the $y$-th entry of $\hat{\mathbf{q}}_{\textsc{IQ}}$ is a weighted average of $\hat{q}$ and $\hat{q}_{y}^{\textsc{CW}}$ (as defined in (4) and (6)) for some weight $\tau\in[0,1]$. That is,
@@
We remark that this theoretical lower bound is likely conservative and, in practice, Interp-Q achieves coverage close to $1-\alpha$.
In Appendix A, we consider an alternative way of instantiating the interpolation idea used by Interp-Q. We propose a method called Fuzzy Classwise CP that interpolates by computing weighted quantiles using class-dependent weights determined by some notion of class similarity.
@@ 3 Experiments
Overview. We consider two datasets for long-tailed classification: Pl@ntNet-300K (Garcin et al., 2021) and iNaturalist-2018 (Van Horn et al., 2018). Figure 2 shows the class distributions of the datasets we use. Unless otherwise stated, we use the $s_{\mathsf{softmax}}$ score function described in Section 1.2. The base model is a ResNet-50 (He et al., 2016) trained using the standard cross-entropy loss (see Appendix D.3 for similar results using focal loss Lin et al., 2017, a loss designed for class-imbalanced settings). More details about the experimental setup are available in Appendix C. Code for reproducing our experiments is available at https://github.com/tiffanyding/long-tail-conformal.
Figure 2: Class distributions (sorted by prevalence), plotted using a logarithmic scale, of the classical train, val, and test sets in the datasets we experiment on. We further randomly split 30% of val to use for model validation and use the remaining 70% as the calibration set ${\mathcal{D}}_{\mathrm{cal}}$. We use the truncated versions when it is important to have good estimates of class-conditional metrics.
@@
as the empirical class-conditional coverage of class $y$. We consider several ways of aggregating $\hat{c}_{y}$ across $y$ to obtain a scalar metric: (i) the fraction of classes with coverage below a threshold (50% in our experiments), (ii) the undercoverage gap, defined as the average undercoverage across classes (zero for classes with coverage at least $1-\alpha$), and (iii) the average of $\hat{c}_{y}$’s, yielding the empirical macrocoverage. This is summarized in the first three rows of Table˜2. Which of these metrics is most natural depends on the goal of the practitioner. Furthermore, we consider the standard prediction set metrics of marginal coverage and average set size, as detailed in the last two rows of Table˜2.
Methods. We run Standard with the $\mathsf{PAS}$ score function from Section 2.1. and Interp-Q from Section˜2.2 with weights $\tau\in$ {0, 0.25, 0.5, 0.75, 0.9, 0.95, 0.975, 0.99 , 0.999, 1} on the Classwise thresholds. These weights are chosen to effectively trace out the trade-off between set size and class-conditional coverage in our experiments. We compare against the following baseline conformal prediction methods: Standard, as described in (4); Classwise, as described in (6); and Clustered (Ding et al., 2023), a method that targets class-conditional coverage in the many-classes setting by grouping together classes with similar score distributions and computing a single score threshold for each cluster.
@@ 3.1 Evaluating the size-coverage trade-off
(i) When targeting set size and class-conditional or macro-coverage, it is more effective to optimize for this trade-off directly than trading off set size and marginal coverage. Adjusting $\alpha$ in Standard CP is a plausible solution to target class-conditional coverage but the results show that this does not optimally navigate the set size/class-conditional coverage trade-off, which is our goal. In comparison, our methods, which explicitly optimize for class-conditional or macro-coverage, consistently achieve better trade-offs than Standard CP.
@@ 3.2 Targeting endangered species
Motivated by plant conservation, we use the weighted prevalence-adjusted softmax $(\mathsf{WPAS})$ score to target coverage of at-risk species in Pl@ntNet-300K. [^1] We consider species with an IUCN status of “endangered”, “vulnerable”, “near threatened”, “critically endangered”, or “lower risk” as at risk. Of the 1,081 total species in Pl@ntNet-300K, 33 species qualify as at-risk. Let ${\mathcal{Y}}_{\text{at-risk}}\subseteq{\mathcal{Y}}$ be the set of at-risk species. The coverage of at-risk species is weighted $\lambda\geq 1$ times more than the coverage of other species, so
@@ 3.3 Simulated human decision-maker
We now examine how coverage and set size can jointly influence human decision accuracy. Human interpretations of prediction sets can vary (Zhang et al., 2024; Hullman et al., 2025), and we focus on two models of human decision-making that are impacted by coverage and set size: an expert verifier $H_{\mathrm{expert}}$ and a random guesser $H_{\mathrm{random}}$ (equivalent to the uncertainty suppressing decision maker in Hullman et al., 2025). Let $H({\mathcal{C}}(X),Y)\in{\mathcal{Y}}$ be a human’s chosen label given prediction set ${\mathcal{C}}(X)$ when the true label is $Y$. The probability $P(H({\mathcal{C}}(X),Y)=Y)$ that the human chooses the correct label after seeing prediction set ${\mathcal{C}}(X)$ is $\mathbbm{1}\{Y\in{\mathcal{C}}(X)\}$ if $H=H_{\mathrm{expert}}$ and $\mathbbm{1}\{Y\in{\mathcal{C}}(X)\}/|{\mathcal{C}}(X)|$ if $H=H_{\mathrm{random}}$. $H_{\mathrm{expert}}$ is only affected by coverage and not set size, whereas random guessers are highly sensitive to set size, as they choose a label uniformly at random from the prediction set. We also consider mixture decision-makers, $H_{\mathrm{mixture}}$, who act as $H_{\mathrm{expert}}$ with probability $\gamma_{\mathrm{exp.}}$ and $H_{\mathrm{random}}$ with $1-\gamma_{\mathrm{exp.}}$ for $\gamma_{\mathrm{exp.}}\in[0,1]$. We are interested in how prediction sets affect the probability that a human correctly labels an instance of a given class, which we formalize as the class-conditional decision accuracy for class $y$ under procedure $\mathcal{C}$}, defined as ${\mathbb{P}}(H({\mathcal{C}}(X),Y)=Y\mid Y=y)$.
Figure 5 shows the class-conditional decision accuracies for various types of decision makers under Standard with $\mathsf{PAS}$ as compared to baselines (the plots for Interp-Q are similar; see Appendix D.6). We report results on truncated versions of Pl@ntNet-300K and iNaturalist-2018 that only include classes with enough test examples to reliably estimate class-conditional decision accuracy (see Appendix C.1 for more details). We observe that Standard and Standard with $\mathsf{PAS}$ achieve strong performance regardless of $\gamma_{\mathrm{exp.}}$, whereas Classwise only does well when $\gamma_{\mathrm{exp.}}$ is high. However, Standard with $\mathsf{PAS}$ has an additional benefit that its performance is more balanced across classes. The worst decision accuracies for Standard with $\mathsf{PAS}$ are higher than those of Standard, which is paid for by only a slight decrease in the best decision accuracies.
Figure 5: Class-conditional decision accuracies for a range of decision makers when presented with sets from Standard, Classwise or Standard with $\mathsf{PAS}$ at $\alpha=0.1$. Classes are ordered by decreasing decision accuracy of $H_{\mathrm{expert}}$ under each method.
@@ 4 Discussion
Limitations and future work. To use Interp-Q, we must choose a value of $\tau$. In practice, if we select a desired average set size and choose the parameter value that achieves this size on the calibration set, this will produce reasonable results (despite the exchangeability violation). More broadly, a promising future research direction that we touched upon is the interaction between set sizes and coverage in determining the utility of prediction sets. In particular, an important aspect of decision-making not addressed by the decision accuracy discussed in Section˜3.3 is the concept of "effort": even if a human can identify the correct label within a set, larger sets require more effort to search through.
@@
Acknowledgments and Disclosure of Funding
We thank Pierre Bonnet, Alexis Joly, Anastasios Angelopolous, and Ryan Tibshirani for insightful conversations and Margaux Zaffran for making this project possible. This work was funded by the French National Research Agency (ANR) through the grant Chaire IA CaMeLOt (ANR-20-CHIA-0001-01).
@@ Appendix A Fuzzy Classwise Conformal Prediction
The Interp-Q method introduced in Section Section˜2.2 shrinks $\hat{q}_{y}$ in a naive way that disregards relationships between classes; in this section, we propose a more sophisticated approach called fuzzy classwise CP (Fuzzy for short) that assigns weights to the score of each calibration example based on the “similarity” between their class and class $y$ and then takes the quantile of this weighted distribution. Notably, this approach comes with a $1-\alpha$ marginal coverage guarantee (rather than the $1-2\alpha$ guarantee of Interp-Q). Empirically, we find that it does not do better than the simpler Interp-Q procedure, but we still believe it is useful to present the method, as it may be preferable in settings where a marginal coverage guarantee is desired. Proofs are deferred to Appendix B.
This method relates to work on weighted conformal prediction, which generalizes standard conformal prediction by computing a weighted quantile of calibration scores (Tibshirani et al., 2019; Barber et al., 2023; Barber and Tibshirani, 2025). Methodologically closest to our work is Podkopaev and Ramdas (2021), which uses label-dependent weighting to ensure marginal coverage under label shift.
@@ A.1 Preliminaries
Recall the Standard CP and Classwise CP procedures introduced in Section˜1.2. We now consider a generalization of these two methods. Let $w:{\mathcal{Y}}\times{\mathcal{Y}}\to{\mathbb{R}}_{\geq 0}$ be a weighting function where $w(y^{\prime},y)$ is the weight assigned to a point with label $y^{\prime}$ when computing the weighted quantile threshold for class $y$. The $w$-label-weighted conformal prediction set (see, e.g., Podkopaev and Ramdas, 2021) is ${\mathcal{C}}_{\mathrm{LW}}(X;w):={\mathcal{C}}(X;\hat{\mathbf{q}}_{w})$ where the $y$-th entry of $\hat{\mathbf{q}}_{w}$ is
@@ A.2 The Fuzzy method
To construct this weighting function, we use a class mapping $\Pi:{\mathcal{Y}}\to\Lambda\subseteq{\mathbb{R}}^{d}$ for some small $d>0$ (e.g., $\Lambda=(0,1)$) and kernel functions $h_{y}^{\sigma}:\Lambda\times\Lambda\to\mathbb{R}_{\geq 0}$ parameterized by a bandwidth $\sigma>0$. A good $\Pi$ should map classes with similar score distributions close together. The kernel $h_{y}^{\sigma}$ takes in two mapped classes and outputs a non-negative scalar that is decreasing with the distance between the two inputs, and it is allowed to depend on $y$. For example, its bandwidth could be rescaled to decrease with the number of examples we have from class $y$ so that classes with more examples “borrow less” from other classes.
@@
Then, we have the equivalence $y\in{\mathcal{C}}_{\textsc{RawFuzzy}}(x)\Longleftrightarrow\tilde{s}(x,y)<1-\alpha$. The score $\tilde{s}$ can be interpreted as (one minus) a weighted conformal p-value (Vovk et al., 2005; Barber and Tibshirani, 2025), which we re-calibrate. To achieve the desired marginal coverage, the method Fuzzy does Standard CP with the new score function $\tilde{s}$ and held-out data as a calibration set.
@@
Note that, if desired, we could use full conformal techniques (Vovk et al., 2005) instead of data splitting (i.e., let ${\mathcal{D}}_{\mathrm{cal}}={\mathcal{D}}_{\mathrm{hold}}$), but this incurs higher computational costs that are undesirable for practical applications (see Proposition˜8 for details).

### v3

@@ 2.2 Approach II: Softening Classwise CP via Interp-Q
In the previous section, we presented our first solution, a new conformal score function; here we present our second solution, which is a simple way to interpolate between the behaviors of Classwise and Standard by linearly interpolating their quantile thresholds. We call this procedure Interp-Q for “interpolated quantile” because it constructs sets as ${\mathcal{C}}_{\textsc{Interp-Q}}(X):={\mathcal{C}}(X;\hat{\mathbf{q}}_{\textsc{IQ}})$ where the $y$-th entry of $\hat{\mathbf{q}}_{\textsc{IQ}}$ is a weighted average of $\hat{q}$ and $\hat{q}_{y}^{\textsc{CW}}$, as defined in (4) and (6), for some weight $\tau\in[0,1]$. That is,
@@
Theoretically, this lower bound is almost tight; we can construct a pathological example where Interp-Q achieves coverage of $1-2\alpha+\alpha^{2}$, which is close to $1-2\alpha$ for small $\alpha$ (see Appendix B.3). Empirically, however, we find that Interp-Q achieves coverage close to $1-\alpha$. This is because the real data we test on do not exhibit the pathologies from our example (which uses discrete score distributions that differ greatly between classes).
In Appendix A, we consider an alternative way of instantiating the interpolation idea used by Interp-Q via a method we call Fuzzy Classwise CP, which interpolates by computing weighted quantiles using class-dependent weights determined by some notion of class similarity.
@@ 3 Experiments
Overview. Code for reproducing our experiments is available at https://github.com/tiffanyding/long-tail-conformal. We consider two datasets for long-tailed classification: Pl@ntNet-300K (Garcin et al., 2021) and iNaturalist-2018 (Van Horn et al., 2018). Figure 2 shows the class distributions of the datasets we use. A key challenge of Pl@ntNet-300K and iNaturalist-201 is that their test sets are also long-tailed, hindering reliable class-conditional evaluation for rare classes.11169% of classes in Pl@ntNet-300K and 90% of classes in iNaturalist have fewer than 10 test examples. To address this, we create truncated versions (see Appendix C.1 for details) that preserve some of the long-tail structure while allowing for robust estimation of class-conditional metrics: Classes in the truncated datasets have 100 test examples each, but we assume the test distribution of interest is equivalent to the (long-tailed) train distribution, so we compute marginal metrics by computing a weighted average of the class-conditional metrics where the weight for class $y$ is the prevalence $\hat{p}(y)$ computed on train. When it comes to aggregated class-conditional metrics (such as macro-coverage), the full and truncated datasets produce similar results, so we defer the truncated results to Appendix D. However, when reporting unaggregated class-conditional metrics (as is the case in Section 3.3 below), it is necessary to use the truncated versions.
Figure 2: Class distributions (sorted by prevalence), plotted using a logarithmic scale, of the classical train, val, and test sets in the datasets we experiment on. We further randomly split 30% of val to use for model validation and use the remaining 70% as the calibration set ${\mathcal{D}}_{\mathrm{cal}}$. We use the truncated versions to obtain good estimates of class-conditional metrics.
Unless otherwise stated, we use the $s_{\mathsf{softmax}}$ score function described in Section 1.2. The base model is a ResNet-50 (He et al., 2016) trained using the standard cross-entropy loss (see Appendix D.3 for similar results using focal loss (Lin et al., 2017), a loss designed for class-imbalanced settings). More details about the experimental setup are available in Appendix C.
as the empirical class-conditional coverage of class $y$.
@@
We consider several ways of aggregating $\hat{c}_{y}$ across $y$ to obtain a scalar metric: (i) the fraction of classes with coverage below a threshold (50% in our experiments), (ii) the undercoverage gap, defined as the average undercoverage across classes (zero for classes with coverage at least $1-\alpha$), and (iii) the average of $\hat{c}_{y}$’s, yielding the empirical macrocoverage. This is summarized in the first three rows of Table 2. Which of these metrics is most natural depends on the goal of the practitioner. Furthermore, we consider the standard prediction set metrics of marginal coverage and average set size, as detailed in the last two rows of Table 2.
Methods. We run Standard with the $\mathsf{PAS}$ score function from Section 2.1. and Interp-Q from Section 2.2 with weights $\tau\in$ {0, 0.25, 0.5, 0.75, 0.9, 0.95, 0.975, 0.99 , 0.999, 1} on the Classwise thresholds. These weights are chosen to effectively trace out the trade-off between set size and class-conditional coverage in our experiments. We compare against the following baseline conformal prediction methods: Standard, as described in (4); Classwise, as described in (6); and Clustered (Ding et al., 2023), a method that targets class-conditional coverage in the many-classes setting by grouping together classes with similar score distributions and computing a single score threshold for each cluster. Additional baselines with weaker performance, such as RC3P from Shi et al. (2024), are deferred to Appendix D.
@@ 3.1 Evaluating the size-coverage trade-off
(i) When targeting set size and class-conditional or macro-coverage, it is more effective to optimize for this trade-off directly than trading off set size and marginal coverage. Adjusting $\alpha$ in Standard CP is a plausible way to target class-conditional coverage but the results show that this does not optimally navigate the set size/class-conditional coverage trade-off, which is our goal. In comparison, our methods, which explicitly optimize for class-conditional or macro-coverage, consistently achieve better trade-offs than Standard CP.
@@ 3.2 Targeting endangered species
Motivated by plant conservation, we use the weighted prevalence-adjusted softmax $(\mathsf{WPAS})$ score to target coverage of at-risk species in Pl@ntNet-300K. [^2] We consider species with an IUCN status of “endangered”, “vulnerable”, “near threatened”, “critically endangered”, or “lower risk” as at risk. Of the 1,081 total species in Pl@ntNet-300K, 33 species qualify as at-risk. Let ${\mathcal{Y}}_{\text{at-risk}}\subseteq{\mathcal{Y}}$ be the set of at-risk species. We will weight the coverage of at-risk species $\lambda\geq 1$ times more than the coverage of other species, so
@@ 3.3 Simulated human decision-maker
We now examine how coverage and set size can jointly influence human decision accuracy. Human interpretations of prediction sets can vary (Zhang et al., 2024; Hullman et al., 2025), and we focus on two models of human decision-making that are impacted by coverage and set size: an expert verifier $H_{\mathrm{expert}}$ and a random guesser $H_{\mathrm{random}}$ (equivalent to the uncertainty suppressing decision maker in Hullman et al., 2025). Let $H({\mathcal{C}}(X),Y)\in{\mathcal{Y}}$ be a human’s chosen label given prediction set ${\mathcal{C}}(X)$ when the true label is $Y$. The probability ${\mathbb{P}}(H({\mathcal{C}}(X),Y)=Y)$ that the human chooses the correct label after seeing prediction set ${\mathcal{C}}(X)$ is $\mathbbm{1}\{Y\in{\mathcal{C}}(X)\}$ if $H=H_{\mathrm{expert}}$ and $\mathbbm{1}\{Y\in{\mathcal{C}}(X)\}/|{\mathcal{C}}(X)|$ if $H=H_{\mathrm{random}}$. $H_{\mathrm{expert}}$ is only affected by coverage and not set size, whereas random guessers are highly sensitive to set size, as they choose a label uniformly at random from the prediction set. We also consider mixture decision-makers, $H_{\mathrm{mixture}}$, who act as $H_{\mathrm{expert}}$ with probability $\gamma_{\mathrm{exp.}}$ and $H_{\mathrm{random}}$ with $1-\gamma_{\mathrm{exp.}}$ for $\gamma_{\mathrm{exp.}}\in[0,1]$. We are interested in how prediction sets affect the probability that a human correctly labels an instance of a given class, which we formalize as the class-conditional decision accuracy for class $y$ under procedure $\mathcal{C}$}, defined as ${\mathbb{P}}(H({\mathcal{C}}(X),Y)=Y\mid Y=y)$. Note that, due to the definition of $H_{\mathrm{expert}}$, the class-conditional decision accuracy of $H_{\mathrm{expert}}$ is equivalent to the class-conditional coverage.
Figure 5 shows the class-conditional decision accuracies for various types of decision makers under Standard with $\mathsf{PAS}$ as compared to baseline methods (the plots for Interp-Q are similar; see Appendix D.6). We report results on the truncated versions of Pl@ntNet-300K and iNaturalist-2018 that only include classes with enough test examples to reliably estimate class-conditional decision accuracy (see Appendix C.1 for more details). We observe that Standard and Standard with $\mathsf{PAS}$ achieve strong performance regardless of $\gamma_{\mathrm{exp.}}$, whereas Classwise only does well when $\gamma_{\mathrm{exp.}}$ is high. However, compared to Standard, Standard with $\mathsf{PAS}$ has an additional benefit that its performance is more balanced across classes. The worst decision accuracies for Standard with $\mathsf{PAS}$ are higher than those of Standard, which is paid for by only a slight decrease in the best decision accuracies.
Figure 5: Class-conditional decision accuracies for a range of decision makers when presented with sets from Standard, Classwise or Standard with $\mathsf{PAS}$ at $\alpha=0.1$ . Classes are ordered by decreasing decision accuracy of $H_{\mathrm{expert}}$ under each method. Note that the decision accuracy when $\gamma_{\mathrm{exp.}}=100\%$ is equivalent to the class-conditional coverage.
@@ 4 Discussion
In this paper, we consider the problem of producing good prediction sets in long-tailed settings, which heretofore has been an understudied problem despite its relevance in many applications, such as species or disease classification. We develop two approaches to better navigate the trade-off between set size and class-conditional coverage in long-tailed settings. First, we introduce prevalence-adjusted softmax ($\mathsf{PAS}$) and its weighted variant ($\mathsf{WPAS}$), score functions inspired by oracle solutions that optimally balance expected set size and (weighted) macro-coverage. Second, we propose Interp-Q, a simple procedure that interpolates between Standard and Classwise CP to provide a tunable compromise between small sets and strong class-conditional coverage.
Limitations and future work. To use Interp-Q, we must choose a value of $\tau$. In practice, if we select a desired average set size and choose the parameter value that achieves this size on the calibration set, this will produce reasonable results (despite the exchangeability violation). Another promising future research direction that we touched upon is the interaction between set sizes and coverage in determining the utility of prediction sets. In particular, an important aspect of decision-making not addressed by the decision accuracy discussed in Section 3.3 is the concept of “effort”: even if a human can identify the correct label within a set, larger sets require more effort to search through.
@@
Acknowledgments
We thank Pierre Bonnet, Alexis Joly, Anastasios Angelopolous, and Ryan Tibshirani for insightful conversations and Margaux Zaffran for making this project possible. This work was funded by the French National Research Agency (ANR) through the grant Chaire IA CaMeLOt (ANR-20-CHIA-0001-01). TD additionally acknowledges support from the National Science Foundation Graduate Research Fellowship Program under grant no. 2146752.
@@ Appendix A Fuzzy Classwise Conformal Prediction
The Interp-Q method introduced in Section 2.2 shrinks $\hat{q}_{y}$ in a naive way that disregards relationships between classes; in this section, we propose a more sophisticated approach called fuzzy classwise CP (Fuzzy for short) that assigns weights to the score of each calibration example based on the “similarity” between their class and class $y$ and then takes the quantile of this weighted distribution. Notably, this approach comes with a $1-\alpha$ marginal coverage guarantee (rather than the $1-2\alpha$ guarantee of Interp-Q). Empirically, we find that it does not do better than the simpler Interp-Q procedure, but we still believe it is useful to present the method, as it may be preferable in settings where a marginal coverage guarantee is desired. Proofs are deferred to Appendix B.
This method relates to work on weighted conformal prediction, which generalizes standard conformal prediction by computing a weighted quantile of calibration scores (Tibshirani et al., 2019; Barber et al., 2023; Barber & Tibshirani, 2025). Methodologically closest to our work is Podkopaev & Ramdas (2021), which uses label-dependent weighting to ensure marginal coverage under label shift.
@@ A.1 Preliminaries
Recall the Standard CP and Classwise CP procedures introduced in Section 1.2. We now consider a generalization of these two methods. Let $w:{\mathcal{Y}}\times{\mathcal{Y}}\to{\mathbb{R}}_{\geq 0}$ be a weighting function where $w(y^{\prime},y)$ is the weight assigned to a point with label $y^{\prime}$ when computing the weighted quantile threshold for class $y$. The $w$-label-weighted conformal prediction set (see, e.g., Podkopaev & Ramdas, 2021) is ${\mathcal{C}}_{\mathrm{LW}}(X;w):={\mathcal{C}}(X;\hat{\mathbf{q}}_{w})$ where the $y$-th entry of $\hat{\mathbf{q}}_{w}$ is
@@ A.2 The Fuzzy method
To construct this weighting function, we use a class mapping $\Pi:{\mathcal{Y}}\to\Lambda\subseteq{\mathbb{R}}^{d}$ for some small $d>0$ (e.g., $\Lambda=(0,1)$) and kernel functions $h_{y}^{\sigma}:\Lambda\times\Lambda\to\mathbb{R}_{\geq 0}$ parameterized by a bandwidth $\sigma>0$. A good $\Pi$ should map classes with similar score distributions close together. The kernel $h_{y}^{\sigma}$ takes in two mapped classes and outputs a non-negative scalar that is decreasing with the distance between the two inputs, and it is allowed to depend on $y$. For example, the kernel bandwidth could be rescaled to decrease with the number of examples we have from class $y$ so that classes with more examples “borrow less” from other classes.
@@
Then, we have the equivalence $y\in{\mathcal{C}}_{\textsc{RawFuzzy}}(x)\Longleftrightarrow\tilde{s}(x,y)<1-\alpha$. The score $\tilde{s}$ can be interpreted as (one minus) a weighted conformal p-value (Vovk et al., 2005; Barber & Tibshirani, 2025), which we re-calibrate. To achieve the desired marginal coverage, the method Fuzzy does Standard CP with the new score function $\tilde{s}$ and held-out data as a calibration set.
@@
Note that, if desired, we could use full conformal techniques (Vovk et al., 2005) instead of data splitting (i.e., let ${\mathcal{D}}_{\mathrm{cal}}={\mathcal{D}}_{\mathrm{hold}}$), but this incurs higher computational costs that are undesirable for practical applications (see Proposition 8 for details).